# ID-Guard: A Universal Framework for Combating Facial Manipulation via Breaking Identification

### v1

@@ IV-C2 Dynamic Weight Strategy
The dynamic weight strategy focuses on balancing the attack losses for different facial manipulations. We selected equivalent weight, prior weight, hard model mining (HMM) [13], and KPI as the baseline of the weight setting methods. The equivalent weight setting will cause the generated perturbations to overfit on the most vulnerable model architecture (e.g., StarGAN and FPGAN). Although HMM balances each model to a certain extent, it ignores the difference in model gradients and thus causes the degradation of average performance. Separate prior weight setting or KPI are unstable and difficult to set, so we cleverly blend the two in Strategy II and get stable training. The benefits of this are two-fold: 1) It reduces the difficulty of a prior setting, and only needs to determine a series of orders of magnitude to allow automatic optimization of parameters; 2) It makes the KPI strategy more stable. Strategy I also achieves excellent results, but the additional backpropagation makes its training more expensive.
@@ IV-C3 Gradient Prior Perturbation
Gradient prior perturbation aims to provide the generator with noise-like prior knowledge, thus accelerating its convergence. For comparison, the variation of training loss and defense performance at different scales with gradient a prior perturbation, with a prior random noise, and without prior knowledge are shown in Fig. 8. Both gradient prior perturbation and random noise promote the convergence of the generator, which is due to the introduction of global noise structure [4]. In terms of generator performance, methods based on gradient prior perturbation at different scales have shown the most significant defense effect, with an average improvement of 31.2% compared to random noise methods. We believe that the reason behind this is that gradient prior perturbation involves rich adversarial structural information.
@@ IV-C4 Architecture of the Perturbation Generator
We explore the impact of different generator architectures on performance. Three mainstream architectures including Unet [52], Resnet [38] and Transformer [53] were selected as the generators of the proposed ID-guard. Fig. IV reports the defense performance of the generators for these three architectures. Compared with Unet, Resnet and Transformer architectures have achieved significant advantages. As shown in Table IV, Transformer achieved optimal performance at the expense of model parameter size, while Resnet achieved very close performance with less than 5% of its parameter size. We propose to use Resnet as the architecture for the generator of the proposed ID-guard, and the intuition behind this is that the generated perturbation can be regarded as a residual of the image.
@@ IV-D1 Misleading Facial Recognition Systems
Some social applications recognize photos uploaded by users and then add corresponding tags and use them in content recommendation systems. This can exacerbate the spread of distorted faces. Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems. Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems. As shown in Fig. 10, we evaluate the misdirection success rates of the destroyed outputs of ID-guard and competing algorithms on three mainstream face recognition systems. As can be seen, our method reports optimal results, achieving over 95% misdirection success rate on Google  [^6] https://www.google.com and StarByFace  [^7] https://starbyface.com. Baidu  [^8] https://www.baidu.com has the most robust recognition system, with CMUA [12] and PG [39] can hardly fool it, but ID-guard still causes it to recognize more than 75% of images incorrectly. The good performance of ID-guard is due to the identity loss introduced to destroy the identity recognition baseline model, which is widely used in commercial face recognition systems.
@@
IV-D2 Resisting Face Inpaintings
Another challenge comes from the image inpainting system. A well-trained face inpainting model can recover distorted facial images, rendering defenses ineffective. We evaluate the performance of distorted images against two baseline face inpainting models, namely LBP [54] and GS-SSA [55]. The quantitative results of $L_{2}^{face}$ distance are reported in Table V, and the visualization results are shown in Fig. 11. Although the difference between the repaired distorted image and the forged result is greatly reduced, the proposed method still exhibits optimal defense performance. This is because these face inpainting systems rely heavily on undistorted regions of the image, However, we achieve a greater degree of destruction of the entire image texture due to the introduction of feature loss.
@@ IV-D2 Resisting Face Inpaintings
  Face inpaintings Methods StarGAN AGGAN FPGAN RelGAN HiSD
@@ IV-D3 Performance in Gray-box Scenarios
The performance of the proposed method in gray-box scenarios is also evaluated. Gray box scenarios are defined where the model type is known but the internal parameters are not accessible. We train a StarGAN [21] and RelGAN [24] as the target gray-box model respectively. As shown in Table VI, to some extent, ID-guard still maintains the most powerful defense capabilities. Furthermore, it can be found that compared to competing algorithms, ID-guard has the smallest performance degradation. We attribute this to the designed feature loss, which destroys model feature extraction with cross-model consistency [17].
@@
IV-D4 Robustness in Real Social Scenarios
@@
In real scenarios, users often upload perturbed images to social applications to share their lives. However, various lossy operations on the transmission channel can destroy the effectiveness of the perturbation. In this section, we first evaluate the robustness of competing algorithms on Facebook and WeChat. Next, we incorporate the robustness strategy in [18] into our generator training, which can be viewed as a downstream task of this work. Quantitative results are shown in Table VII. Even without robustness training, our method still reports the best robustness compared to the baseline. The reason may be that our adversarial perturbation is not averaged over the image, but is constrained to focus on faces, which makes it harder to neutralize. As shown in Fig. 12, when the robust training strategy is integrated, the anti-lossy operation performance of generated perturbations is greatly improved. This demonstrates the flexibility of the proposed ID-guard framework to effectively integrate with progressive strategies in the community.
TABLE VII: Quantitative results on the robustness of adversarial perturbations in real scenarios. The adversarial images are uploaded to designated social applications and then downloaded. The best result is marked in bold. “w/” and “w/o” indicate whether robust training is used or not, respectively.
[…]
Facebook Disrupting[10] 0.023 0.760 0.022
PG[39] 0.012 0.744 0.009
CMUA[12] 0.003 0.761 0.001
IAP[4] 0.012 0.749 0.016
Ours (S-II) w/o 0.041 0.708 0.112
Ours (S-II) w/ 0.065 0.677 0.306
  WeChat Disrupting[10] 0.010 0.762 0.006
PG[39] 0.004 0.747 0.003
CMUA[12] 0.002 0.741 0.001
IAP[4] 0.004 0.746 0.001
Ours (S-II) w/o 0.017 0.739 0.024
Ours (S-II) w/ 0.052 0.702 0.134
Figure 12: Visual examples of defense effects in real scenarios. The transmission channel is Facebook and the StarGAN is chosen as the target model. “w/” and “w/o” indicate whether robust training is used or not, respectively.
@@ IV-E1 How Weights Dynamically Change?
In the proposed ID-guard framework, the dynamic weight strategy is very important, which directly affects the training process and the balance of attack losses for different facial manipulations. Here, to explore its mechanism in depth, we record the dynamic changes of the weight set $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ of attack losses in an epoch training, as shown in Fig. 13. For strategy I, we use the MGDA algorithm to solve the set of weights automatically in each iteration. It can be found that face manipulation models with strong robustness, such as HiSD [25] and RelGAN [24], tend to be assigned larger weights to strengthen the attack against them. In addition, due to the lack of prior knowledge guidance, the weight change of strategy I is more affected by the current state of the generator, and thus fluctuates more significantly. For strategy II, in the initial stage, there is a large difference between the weights. With the constraints of the prior weights, the allocation of each dynamic weight stabilizes to obtain a balanced performance.
@@
IV-E2 What does the generator learn?
The adversarial perturbations generated by the generator trained under different loss constraints are shown in Fig. 14. The mask loss concentrates rich adversarial information on the facial area of the image, thereby completely distorting the output face. The generator trained using only the identity loss learns to disrupt images at the texture level. Combined with the results in Fig 7, this destruction changes the key feature semantics and visual attributes of the face. Therefore, the standard generator learns to generate adversarial perturbations that cause maximum damage to facial regions and alter the identifiable texture features of the image.
Figure 14: Visual examples of the generated perturbation. Among them, (a) is the original image, (b) is the adversarial image, (c) is the perturbation generated by the generator trained only with mask loss, (d) is the perturbation generated by the generator trained only with identity loss, and (e) is the perturbation generated by the standard generator.
@@ V Conclusion
In this work, we proposed a universal adversarial framework for combating facial manipulation, named ID-guard. To prevent face stigmatization problems caused by unconstrained image distortion, we propose an identity destruction module to eliminate identity semantics. Furthermore, to improve the cross-model performance of generating perturbations, we regard attacking different models as a multi-task learning problem and introduce a dynamic parameter strategy. The proposed method not only effectively resists multiple face manipulations, but also significantly disrupts face identification. In addition, the experiment also demonstrated the possibility of ID-guard in circumventing commercial face recognition systems and image inpaintings. We hope that ID-guard, with its good integration capabilities and application flexibility, can provide the community with an effective solution against facial manipulation.

### v2

@@ IV-C1 Identity Destruction Module
          FCL means the Feature Confusion Loss.
          ICL means the Identity Consistency Loss.
          MCL means the Mask Constrained Loss.
@@ IV-C2 Dynamic Weight Strategy
The dynamic weight strategy focuses on balancing the attack losses for different facial manipulations. We selected equivalent weight, prior weight, hard model mining (HMM) [24], and KPI as the baseline of the weight setting methods. The equivalent weight setting will cause the generated perturbations to overfit on the most vulnerable model architecture (e.g., StarGAN and FPGAN). Although HMM balances each model to a certain extent, it ignores the difference in model gradients and thus causes the degradation of average performance. Separate prior weight settings or KPI are unstable and difficult to set, so we cleverly blend the two in Strategy II and get stable training. The benefits of this are two-fold: 1) It reduces the difficulty of a prior setting, and only needs to determine a series of orders of magnitude to allow automatic optimization of parameters; 2) It makes the KPI strategy more stable. Strategy I also achieves excellent results, but the additional backpropagation makes its training more expensive.
@@ IV-C3 Gradient Prior Perturbation
Gradient prior perturbation aims to provide the generator with noise-like prior knowledge, thus accelerating its convergence. For comparison, the variation of training loss and defense performance at different scales with the gradient prior perturbation, with a prior random noise, and without prior knowledge is shown in Fig. 9. Both gradient prior perturbation and random noise promote the convergence of the generator, which is due to the introduction of global noise structure [9]. In terms of generator performance, methods based on gradient prior perturbation at different scales have shown the most significant defense effect, with an average improvement of $31.2\%$ compared to random noise methods. We believe that the reason behind this is that the gradient prior perturbation involves rich adversarial structural information.
@@ IV-C4 Architecture of the Perturbation Generator
We explore the impact of different generator architectures on performance. Three mainstream architectures, including Unet [43], Resnet [51], and Transformer [55], are selected as the generators of the proposed ID-Guard. Fig. IV reports the defense performance of the generators for these three architectures. Compared with Unet, Resnet and Transformer architectures have achieved significant advantages. As shown in Table IV, Transformer achieved optimal performance at the expense of model parameter size, while Resnet achieved very close performance with less than $5\%$ of its parameter size. We propose to use Resnet as the architecture for the generator of the proposed ID-Guard, and the intuition behind this is that the generated perturbation can be regarded as a residual of the image.
@@ IV-D1 Misleading Facial Recognition Systems
Some social applications recognize photos uploaded by users and then add corresponding tags and use them in content recommendation systems. This can exacerbate the spread of distorted faces. Therefore, the threat of stigmatization of distorted images comes not only from the human eye but also from commercial facial recognition systems. As shown in Fig. 11, we evaluate the misdirection success rates of the destroyed outputs of ID-Guard and competing algorithms on three mainstream face recognition systems. As can be seen, our method reports optimal results, achieving over $95\%$ misdirection success rate on Google  [^6] https://www.google.com and StarByFace  [^7] https://starbyface.com. Baidu  [^8] https://www.baidu.com has the most robust recognition system, with CMUA [11] and PG [10] can hardly fool it, but ID-Guard still causes it to recognize more than $75\%$ of images incorrectly. The good performance of ID-Guard is due to the identity loss introduced to destroy the identity recognition baseline model, which is widely used in commercial face recognition systems.
@@
IV-D2 Resisting facial inpainting
Another challenge comes from the image inpainting system. A well-trained facial inpainting model can recover distorted facial images, rendering defenses ineffective. We evaluate the performance of distorted images against two baseline facial inpainting models, namely LBP [53] and GS-SSA [54]. The quantitative results of $L_{2}^{face}$ distance are reported in Table V, and the visualization results are shown in Fig. 12. Although the difference between the repaired distorted image and the forged result is greatly reduced, the proposed method still exhibits optimal defense performance. This is because these facial inpainting systems rely heavily on undistorted regions of the image, However, we achieve a greater degree of destruction of the entire image texture due to the introduction of feature loss.
@@ IV-D2 Resisting facial inpainting
facial inpainting Methods StarGAN AGGAN FPGAN RelGAN HiSD
@@ IV-D3 Performance in Gray-box Scenarios
The performance of the proposed method in gray-box scenarios is also evaluated. Gray box scenarios are defined where the model type is known but the internal parameters are not accessible. We train a StarGAN [1] and RelGAN [4] as the target gray-box model, respectively. As shown in Table VI, Disrupting derives adversarial perturbations on the accessible white-box versions of StarGAN and RelGAN, respectively. This model-specific perturbation generation enables it to maintain gray-box transferability to a certain extent. ID-Guard still maintains the most powerful defense capabilities. We attribute this to the designed feature loss, which destroys model feature extraction with cross-model consistency [25]. Additionally, the baseline methods perform better against the gray-box RelGAN than against its white-box version. The reason for this could be that their imbalanced training process caused the perturbation performance to be biased toward the most vulnerable white-box StarGAN, which has adversarial gradients that are closer to the gray-box RelGAN used in this experiment.
@@
IV-D4 Robustness under Lossy Operations
@@
In real scenarios, users often upload perturbed images to social applications to share their lives. However, various lossy operations on the transmission channel can destroy the effectiveness of the perturbation. In this section, we evaluate the robustness of ID-Guard and competing algorithms under JPEG compression and Gaussian blur. To further verify the integration capability of our framework and the robustness strategy, for JPEG, we incorporate the compression-resistant strategy from [27] into our generator training. The results are reported as “Ours (S-II) w/ R.”. As shown in Fig. 13, as the intensity of the lossy operation increases, the defense performance of each method is gradually weakened. Our method demonstrates significant robustness at different scales. The underlying reason for this could be the introduction of the identity disruption module, which concentrates the effectiveness of the adversarial perturbations in specific areas, making them less susceptible to degradation from lossy operations. When the robustness training strategy was integrated, ID-Guard’s robustness improved significantly. A sample visualization is shown in Fig. 14. This demonstrates the flexibility of the proposed framework and its ability to effectively integrate with advanced strategies from the research community.
TABLE VII: Quantitative results of the adversarial robustness of the original face manipulation model and its adversarial training version against different attacks.
Model Gaussian noise ($\sigma=0.05$) FGSM I-FGSM PGD C&W
$L_{2}^{face}$downarrow ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$ $L_{2}^{face}$$\downarrow$ ID sim.$\uparrow$
StarGAN 0.070 0.261 0.185 0.184 1.075 0.069 1.102 0.004 1.377 0.009
StarGAN-AT 0.001 0.985 0.011 0.977 0.148 0.700 0.141 0.712 0.149 0.756
RelGAN 0.003 0.905 0.140 0.766 0.584 0.186 0.615 0.145 0.983 0.099
RelGAN-AT 0.000 0.992 0.001 0.989 0.019 0.783 0.018 0.791 0.002 0.988
(a) JPEG compression
(b) Gaussian Blur
Figure 13: The performance of the defense algorithm against StarGAN under different intensities of lossy operations. In the JPEG evaluation, the performance of ID-Guard integrated with the compression-resistant strategy has a significant advantage. Therefore, in order to present the results, we take the logarithm of 10 for $L_{2}^{face}$.
Figure 14: Visualization of adversarial examples against the StarGAN model under JPEG compression with QF=85.
IV-D5 ID-Guard for Adversarial Training
In this section, we evaluate the effectiveness of using the ID-Guard framework for adversarial training of facial manipulation models. Specifically, we select StarGAN[1] and RelGAN[4] as example facial manipulation models and apply the adversarial training (AT) approach proposed in Section III-E to obtain StarGAN-AT and RelGAN-AT. The adversarial training version of the facial manipulation model $\mathcal{M}$ is noted as $\mathcal{M}^{AT}$. The training framework follows their official open-source implementations  [^9] https://github.com/yunjey/stargan, https://github.com/elvisyjlin/RelGAN-PyTorch. To evaluate adversarial robustness, we test these models against Gaussian noise ($\sigma=0.05$), FGSM[36], I-FGSM[66], PGD[37], and C&W[65] attacks. All attacks follow the white-box setup. As reported in Table VII, non-adversarial Gaussian noise and the single-step FGSM disrupt $\mathcal{M}$ to some extent but have almost no impact on $\mathcal{M}^{AT}$. Furthermore, for the three stronger attack algorithms, I-FGSM, PGD, and C&W, the model implemented adversarial training demonstrates significant robustness. Especially for RelGAN-AT, the $L_{2}^{face}$ consistently remains within the threshold of $0.05$. Additionally, the quantitative identity similarity results and the visualized examples in Fig. 15 indicate that these adversarial attacks mainly distort the background regions of $\mathcal{M}$’s manipulated outputs while leaving the facial regions unaffected. This validates that the proposed ID-Guard framework can serve as a plug-and-play adversarial attack module within adversarial training, significantly enhancing the robustness of facial manipulation models.
Figure 15: Visual example of the adversarial robustness of the original face manipulation model and its adversarial training version against PGD[37].
@@ IV-E1 How Weights Dynamically Change?
In the proposed ID-Guard framework, the dynamic weight strategy is very important, which directly affects the training process and the balance of attack losses for different facial manipulations. Here, to explore its mechanism in depth, we record the dynamic changes of the weight set $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ of attack losses in an epoch training, as shown in Fig. 16. For strategy I, we leverage the MGDA algorithm to optimize the set of weights automatically in each iteration. It can be found that face manipulation models with strong robustness, such as HiSD [5] and RelGAN [4], tend to be assigned larger weights to strengthen the attack against them. In addition, due to the lack of prior knowledge guidance, the weight change of strategy I is more affected by the current state of the generator, and thus fluctuates more significantly. For strategy II, in the initial stage, there is a large difference between the weights. With the constraints of the prior weights, the allocation of each dynamic weight stabilizes to obtain a balanced performance.
@@
IV-E2 Can ID-guard be migrated to active forensics?
We migrate the proposed ID-Guard framework to proactive forensics for facial manipulation to demonstrate its scalability. Specifically, we follow the setup in [14]: Given a natural image $x$ and a binary message $m\in\{0,1\}^{L}$ of length $L$, a message embedding encoder $E_{m}$ maps the image and a corresponding message to a watermarked image. After the watermarked image undergoes manipulation, a message extractor decoder $D_{m}$ is used to recover $m$ from the forged image. The training objective is to minimize $\sum_{k=1}^{N}L^{wm}_{k}=MSE(m,D_{m}(\mathcal{M}_{k}(E_{m}(x,m))))$, while an additional discriminator is employed to ensure the visual quality of the watermarked image. FaceTagger[59] and information-based IAP[14] are selected as baselines. In our method, the proposed dynamic weight strategy is applied to weight $L^{wm}_{k}$ for different $\mathcal{M}_{k}$ during training. The Bit Error Rate (BER) and the visual quality of the watermarked image are quantitatively reported in Table VIII. FaceTagger is trained with StarGAN[1] as the target model, leading to overfitting and consequently poor overall performance. Compared to IAP, our method demonstrates more balanced forensic performance across different facial manipulation models, particularly against HiSD[5], which exhibits significant structural differences. This result highlights the strong scalability and transferability of the proposed framework, making it a plug-and-play tool adaptable to various cross-model tasks in the community.
TABLE VIII: Quantitative results of proactive forensics. The best result is marked in bold.
[…]
StarGAN AGGAN FPGAN RelGAN HiSD
FaceTagger[59] 0.000 0.501 0.009 0.069 0.512 0.995
IAP[14] 0.034 0.144 0.000 0.001 0.000 0.983
Ours 0.000 0.000 0.000 0.000 0.000 0.995
IV-E3 What does the generator learn?
The adversarial perturbations generated by the generator trained under different loss constraints are shown in Fig. 17. The mask loss concentrates rich adversarial information on the facial area of the image, thus completely distorting the output face. The generator trained using only the identity loss learns to disrupt images at the texture level. Combined with the results in Fig. 8, this destruction changes the key feature semantics and visual attributes of the face. Therefore, the standard generator learns to generate adversarial perturbations that cause maximum damage to facial regions and alter the identifiable texture features of the image.
Figure 17: Visual examples of the generated perturbation. Among them, (a) is the original image, (b) is the adversarial image, (c) is the perturbation generated by the generator trained only with mask constrained loss, (d) is the perturbation generated by the generator trained only with identity consistency loss, and (e) is the perturbation generated by the standard generator.
@@ V Conclusion
In this work, we proposed a universal framework for combating facial manipulation, named ID-Guard. To prevent face stigmatization problems caused by unconstrained image distortion, we propose an Identity Destruction Module to eliminate identity semantics. Furthermore, to improve the cross-model performance of generating perturbations, we regard attacking different models as a multi-task learning problem and introduce a dynamic parameter strategy. The proposed method not only effectively resists multiple face manipulations but also significantly disrupts face identification. In addition, the experiment also demonstrated the possibility of ID-Guard in circumventing commercial face recognition systems and image inpaintings. We hope that ID-Guard, with its good integration capabilities and application flexibility, can provide the community with an effective solution against facial manipulation.